OpenStudy (ttop0816):

Solve the inequality -9x/11 - 7 < -5. x < -22/9 x > -22/9 x < 22/9 x > 22/9 <--- would this be correct?

OpenStudy (ttop0816):

@kelliegirl33 please help! :D

OpenStudy (mertsj):

1. Add 7 to both sides. 2. Multiply both sides by - 11/9 and don't forget to reverse the inequality symbol.

OpenStudy (mertsj):

\[2\times \frac{-11}{9}=\frac{-22}{9}\]

OpenStudy (ttop0816):

@Mertsj then it would be the second answer from the top?

OpenStudy (mertsj):

yes

OpenStudy (ttop0816):

Solve the inequality 3x - 5 < 4. x < 1/3 x > 1/3 x < 3 <--- it would be this one right?? x > 3

OpenStudy (mertsj):

yes